# The Mechanics of Nucleation and Growth and the Surface Tensions of Active Matter

### v1

@@ II.1 Equilibrium Nucleation Theory
We briefly recapitulate the thermodynamic arguments of CNT applied to bubble/droplet nucleation, the focus of this work. The complete derivation can be found in Section 3 of the Supporting Information (SI)  [^1] See Supporting Information at [URL]. We consider a macroscopic homogeneous single-component system prepared with a uniform number density $\rho^{\rm sat}$ within a fixed system volume $V$ (the total number of particles is thus also fixed at $N=V\rho^{\rm sat}$). We focus on saturation densities $\rho^{\rm sat}$ within the liquid-gas binodal of the phase diagram but far from the spinodal boundaries, limiting our analysis to be far from the critical point. In this region of the phase diagram, a system prepared with a spatially uniform density of $\rho^{\rm sat}$ will remain linearly stable against density fluctuations until a critical nucleus with density $\rho^{\rm in}$ and volume $V^{\rm nuc}$ is formed. We assume that the formation of the critical nucleus is a sufficiently rare event that we may consider a single nucleation event in isolation and a nucleus volume, $V^{\rm nuc}$, that is much smaller than the system volume. These assumptions provide additional limits on the range of $\rho^{\rm sat}$ in which CNT is applicable.
@@
We can make use of the condition that the droplet volume is much smaller than the system volume such that $\rho^{\rm out}\approx\rho^{\rm sat}$. In this limit, we have effectively changed our thermodynamic ensemble as the nucleus is now in contact with a particle reservoir with fixed chemical potential $\mu(\rho^{\rm sat})\equiv\mu^{\rm sat}$ and thus $\rho^{\rm in}$ can be found directly from $\mu(\rho^{\rm in})=\mu^{\rm sat}$. The critical radius is then simply ${R_{C}=2\upgamma/(\mu^{\rm sat}\Delta\rho-\Delta f})$ where the difference between an arbitrary property, $a$, in the interior and exterior of the nucleus is defined as $\Delta a\equiv a^{\rm in}-a^{\rm sat}$. The free energy barrier associated with the critical nucleus can be expressed as the sum of the volumetric and interfacial contributions ${F^{\rm{\dagger}}=4\pi R_{C}^{2}\upgamma-\frac{4}{3}\pi R_{C}^{3}(\mu^{\rm sat}\Delta\rho-\Delta f)}$ or equivalently, ${F^{\rm{\dagger}}=\frac{4}{3}\pi\upgamma R_{C}^{2}}$ [33]. The probability of observing such a critical nucleus is thus expected to be $P(R_{C})\sim\text{exp}\left[F^{{\dagger}}/k_{B}T\right]$.
It is clear that in passive systems, thermodynamics alone provides a guide for determining the density, size, and statistics of the critical nucleus. Given a saturation density, Eqs. (1) are the two equations that allow for the determination of the two unknowns: $\rho^{\rm in}$ and $R_{C}$. In the following section, we derive analogous equations capable of determining $\rho^{\rm in}$ and $R_{C}$ for systems arbitrarily far from equilibrium by adopting a mechanical perspective.
@@ II.2 Mechanics of a Critical Nucleus
Let us consider the dynamical definition of a critical nucleus. When the critical nucleus is an unstable fixed point  [^2] Recently Tjhung et al. [23] demonstrated in some active systems the critical nucleus may be a stable fixed point, in such a case nuclei smaller (larger) than $R_{C}$ will grow (shrink)., nuclei smaller (larger) than $R_{C}$ will be associated with a particle flux normal to the interface such that the nuclei shrink (grow). $R_{C}$ thus represents a critical size at which the direction of this flux switches, so the system must have zero flux normal to the interface precisely when a nucleus has a radius of $R_{C}$. For a single realization of an active fluid, the stochastic noise plays a crucial role in the density flux, which we will discuss further in the next section. For the purposes of determining the properties of the critical nucleus, we can consider an ensemble of supersaturated states. We choose a spherical coordinate system with an origin coinciding with that of the nucleus. The ensemble-averaged nucleus shape will be spherical and the ensemble-averaged density flux will no longer be stochastic. We will determine the critical nucleus size by searching for states of zero radial flux.
@@
Two flux-free states are well-established: the macroscopically phase separated state that the system will eventually reach and the initial metastable homogeneous state. The critical nucleus will correspond to an additional state in which a finite-size minority phase with zero radial flux coexists with a parent phase. The condition for determining a stationary state of a density profile is given by the linear momentum balance (see Section 1 of the SI [33] and Ref. [35]) which governs the dynamics of the density flux. On timescales where inertia can be neglected (or for all times for overdamped particle dynamics), the radial-component of the linear momentum balance is simply $\left(\bm{\nabla}\cdot\bm{\sigma}+\mathbf{b}\right)\cdot\mathbf{e}_{r}=0$, where $\bm{\sigma}$ is the stress tensor, $\mathbf{b}$ are the body forces, and $\mathbf{e}_{r}$ is the unit vector in the radial dimension. We note that exact microscopic expressions for both $\bm{\sigma}$ and $\mathbf{b}$ can be determined through an Irving-Kirkwood-Noll procedure [36, 37, 38], an approach that we will later adopt in applying this theory to active Brownian particles. In these formulations, terms often appear to be non-local, requiring integrals over all space to, for instance, compute the stress at a particular point in space arising from particle interactions. Nevertheless, these terms are often well-approximated (e.g., when interactions are short-ranged and partial spatial correlations decay quickly) by a series of local terms, the coefficients of which can be systematically determined [39]. We thus focus on systems in which $\bm{\sigma}$ and $\mathbf{b}$ can be fully expressed as local density gradient expansions.
@@
Up to a divergence-free term, we may absorb the body force into a stress-like contribution such that ${\bm{\nabla}\cdot\bm{\sigma}+\mathbf{b}=\bm{\nabla}\cdot\left[\bm{\sigma}+\mathbf{F}\right]=\bm{\nabla}\cdot\bm{\Sigma}}$, where $\mathbf{F}$ is the Coulomb integral of $\mathbf{b}$ and $\bm{\Sigma}\equiv\bm{\sigma}+\mathbf{F}$ is the dynamic stress tensor. While $\bm{\sigma}$ and $\mathbf{b}$ are fully local, the Coulomb integral may result in non-local contributions to $\mathbf{F}$. The momentum balance (${\bm{\nabla}\cdot\bm{\Sigma}=\bm{0}}$) can be used to solve for the density profile upon identifying the precise form of the dynamic stress tensor. Under spherical symmetry, $\bm{\Sigma}$ has only two independent components and can be expressed as $\bm{\Sigma}=\Sigma_{rr}\mathbf{e}_{r}\mathbf{e}_{r}+\Sigma_{tt}\left(\mathbf{I}-\mathbf{e}_{r}\mathbf{e}_{r}\right)$ where $\mathbf{I}$ is the identity tensor. We can make analytical progress by expanding the dynamic stress tensor to second order in density gradients while leaving any potential non-local contributions symbolic:
$\Sigma_{rr}=$ $-\mathcal{P}(\rho)+\kappa_{1}(\rho)\left(\frac{\partial}{\partial r}+\frac{2}{r}\right)\frac{\partial\rho}{\partial r}$
$+\bigl{(}\kappa_{2}(\rho)+\kappa_{3}(\rho)\bigr{)}\left(\frac{\partial\rho}{\partial r}\right)^{2}+\kappa_{4}(\rho)\frac{\partial^{2}\rho}{\partial r^{2}}+\mathcal{K}_{r},$ (2a)
$\Sigma_{tt}=-\mathcal{P}(\rho)+\kappa_{1}(\rho)\left(\frac{\partial}{\partial r}+\frac{2}{r}\right)\frac{\partial\rho}{\partial r}+\kappa_{3}(\rho)\left(\frac{\partial\rho}{\partial r}\right)^{2}+\mathcal{K}_{t},$ (2b)
@@
where $-\mathcal{P}(\rho)$ is the bulk isotropic contribution to the dynamic stress tensor, $\mathcal{K}_{r/t}$ are the non-local terms originating from the Coulomb integral of $\mathbf{b}$, and $\{\kappa_{i}\}$ are interfacial coefficients. These interfacial coefficients can be found by formally carrying out the expansion of $\bm{\sigma}$ and $\mathbf{b}$ beginning from the microscopic definition of stress and expanding the local terms that arise from the Coulomb integral of $\mathbf{b}$. The procedure for determining the precise forms of $\{\kappa_{i}\}$ and $\mathcal{K}_{r/t}$ from the expansion coefficients of $\bm{\sigma}$ and $\mathbf{b}$ is outlined in Section 1 of the SI [33]. In a passive system $\mathcal{P}(\rho)$ is simply the bulk pressure and $\{\kappa_{i}\}$ are given by the Korteweg expansion [40] while non-local terms can be well-approximated by local terms for single-component systems with short-ranged interactions. More generally, the bulk pressure and interfacial coefficients can contain dynamical contributions otherwise absent in passive systems [26].
@@
Given the above physical picture, we now outline our process to find the zero radial flux states associated with a finite-sized minority phase [33]. Our spherically symmetric geometry results in the linear momentum balance simplifying to:
@@
$\Delta\mathcal{P}=\Delta\mathcal{K}_{r}+\frac{2\upgamma_{\rm mech}}{R_{C}}+\mathcal{O}\left(\frac{1}{R_{C}^{2}}\right).$ (4a)
@@
$=-\int_{0}^{\infty}\left[\kappa_{3}(\rho)\left(\frac{\partial\rho}{\partial r}\right)^{2}+\kappa_{4}(\rho)\frac{\partial^{2}\rho}{\partial r^{2}}+\mathcal{K}_{r}-\mathcal{K}_{t}\right]\text{d}r,$ (4b)
@@
where the second equality utilizes our specific form of the dynamic stress [see Eqs. (2)]. Equation (4a) is a generalized Young-Laplace equation – with the equilibrium surface energy and pressure more generally replaced by the mechanical tension and dynamic pressure, respectively, and an additional contribution arising from the non-local terms. This mechanical tension was determined to be strongly negative for phase-separated ABPs [21] and has been the subject of some controversy regarding its physical relevance. Solon et al. [30] correctly identified that the mechanical surface tension of Kirkwood and Buff controls the Laplace pressure difference of droplets consistent with a fully local (dynamic) stress. More generally, $\Delta\mathcal{K}_{r}$ may also contribute to the Laplace pressure difference.
@@
While the generalized Young-Laplace equation provides a mechanical condition remarkably similar to its equilibrium counterpart, we still require a second condition to allow us to determine our two unknowns, $\rho^{\rm in}$ and $R_{C}$. For passive fluids, this condition is equality of chemical potential between the nucleus and the surrounding phase. The search for chemical potential-like quantities for nonequilibrium systems has been the subject of much interest in the last decade in the context of determining macroscopic binodals [19, 42, 20, 35]. These works have resulted in generalized Maxwell constructions [20, 35] that are derived from performing the appropriate integration of the mechanical criterion of coexistence [43, 35].
@@
We follow similar arguments used in constructing the nonequilibrium coexistence criteria of macroscopic binodals [43, 35]. Beginning from the mechanical balance [Eq. (3)] we perform an indefinite integral to find an integro-differential equation for $\Sigma_{rr}$ with the Laplace pressure used as a boundary condition. We then introduce the generalized Maxwell construction pseudovariable [43, 20, 35] $\mathcal{E}$ defined by the differential equation ${(\kappa_{1}+\kappa_{4})\partial^{2}\mathcal{E}/\partial\rho^{2}=(2(\kappa_{2}+\kappa_{3})-\partial(\kappa_{1}+\kappa_{4})/\partial\rho)\partial\mathcal{E}/\partial\rho}$, which is found to be $\mathcal{E}\sim 1/\rho$ for a passive system. We multiply the indefinite integral of Eq. (3) by $d\mathcal{E}/dr$ and integrate from $r=0$ to $r=\infty$ to arrive at:
$\int_{\mathcal{E}^{\rm in}}^{\mathcal{E}^{\rm sat}}\text{d}\mathcal{E}$ $\left[\mathcal{P}(\rho)-\mathcal{P}(\rho^{\rm sat})\right]=\frac{2\upgamma_{\rm ost}}{R_{C}}\frac{\rho^{\rm sat}\Delta\mathcal{E}}{\Delta\rho}+\mathcal{O}\left(\frac{1}{R_{C}^{2}}\right),$ (5a)
@@
$\upgamma_{\rm ost}=\frac{\Delta\rho}{\rho^{\rm sat}\Delta\mathcal{E}}$ $\Biggl{[}\int_{0}^{\infty}\Bigl{(}\kappa_{1}(\rho)\frac{\partial\mathcal{E}}{\partial\rho}\left(\frac{\partial\rho}{\partial r}\right)^{2}$
$+\mathcal{G}$ $\left(\mathcal{E}-\mathcal{E}^{\rm in}\right)+\frac{R_{C}}{2}\frac{\partial\mathcal{K}_{r}}{\partial r}\left(\mathcal{E}^{\rm in}-\mathcal{E}\right)\Bigr{)}\text{d}r\Biggr{]}.$ (5b)
Note that $\upgamma_{\rm ost}$ has been defined such that it has the same units as mechanical surface tension and becomes identical to $\upgamma_{\rm mech}$ in a passive system [33].
@@
The connection between Eqs. (4a) and (5a) to the thermodynamic criteria shown in Eqs. (1) can be made explicit. As described by Solon et al., the generalized Maxwell constructions on the pressure can also be expressed as a difference in chemical potential-like quantities between coexisting phases [30]. It was subsequently shown that this potential satisfies a generalized Gibbs-Duhem relation [44], which, for the single-component system under consideration, is expressed as $\text{d}g=\mathcal{E}\text{d}\mathcal{P}$ where $g$ is the generalized chemical potential. The generalized Gibbs-Duhem relation allows [33] Eq. (5a) to be expressed as a difference of chemical potentials with the nonequilibrium critical nucleus criteria now taking the form:
@@
$\Delta g=\mathcal{E}^{\rm in}\Delta\mathcal{K}_{r}+\frac{2}{R_{C}}\left[\mathcal{E}^{\rm in}\upgamma_{\rm mech}-\frac{\rho^{\rm sat}\Delta\mathcal{E}}{\Delta\rho}\upgamma_{\rm ost}\right],$ (6a)
$\Delta\mathcal{P}=\Delta\mathcal{K}_{r}+\frac{2\upgamma_{\rm mech}}{R_{C}}.$ (6b)
@@
Additionally, by defining a pseudopotential ${\Phi\equiv\int\mathcal{P}\text{d}\mathcal{E}}$, Eq. (5a) can be rearranged to express the critical radius as ${R_{C}=2\upgamma_{\rm ost}\rho^{\rm sat}\Delta\mathcal{E}/\Delta\rho\left(\Delta\mathcal{E}\mathcal{P}^{\rm sat}-\Delta\Phi\right)}$. For passive systems without non-local terms, $\upgamma_{\rm mech}=\upgamma_{\rm ost}$ and $\mathcal{E}\sim 1/\rho$. One can use these relations to readily verify that $g^{\rm in}=g^{\rm sat}$ and that Eq. (5a) recovers the thermodynamically derived expression for $R_{C}$ [33]. Our mechanical criteria for the critical nucleus [Eqs. (6)] thus recover and generalize the thermodynamic criteria [Eq. (1)].
Equations (6) [or, equivalently, Eqs. (4) and (5)] are the finite-size coexistence criteria for the critical nucleus. In Ref. [23], similar finite-sized coexistence criteria were derived with two key differences. First, the Laplace pressure difference in Ref. [23] was found to be controlled by the tension controlling Ostwald ripening rather than $\upgamma_{\rm mech}$. Additionally, the chemical potentials (including non-local terms) inside and outside of the nucleus were found to be equal by Ref. [23]. Fundamentally, this is because pressure and chemical potential have different physical interpretations in Ref. [23] than in this work. Here, pressure is the bulk isotropic component of the dynamic stress tensor and carries the familiar mechanical interpretation, while in Ref. [23] the pressure (also referred to as a “pseudopressure”) cannot be physically interpreted without identifying the pseudovariable. In Ref. [23] the chemical potential is defined as the scalar field controlling the dynamics and carries much of the familiar intuition of a chemical potential, while $g$ in this work cannot be physically interpreted without identifying $\mathcal{E}$. The mechanical approach presented above may be a more natural starting point when beginning with microscopic equations-of-motion and using an Irving-Kirkwood approach to derive the form of the dynamic stress and offers the advantage of clearly identifying the role of the mechanical tension of Kirkwood and Buff.

### v2

@@ II.1 Equilibrium Nucleation Theory
We briefly recapitulate the thermodynamic arguments of CNT applied to bubble/droplet nucleation, the focus of this work. The complete derivation can be found in Section 3 of the Supplementary Material (SM). We consider a macroscopic homogeneous single-component system prepared with a uniform number density $\rho^{\rm sat}$ within a fixed system volume $V$ (the total number of particles is thus also fixed at $N=V\rho^{\rm sat}$). We focus on saturation densities $\rho^{\rm sat}$ within the liquid-gas binodal of the phase diagram but far from the spinodal boundaries, limiting our analysis to be far from the critical point. This limitation is necessary in order to ensure that the pathway of the phase transition is unambiguously nucleation and growth, rather than spinodal decomposition. In this region of the phase diagram, a system prepared with a spatially uniform density of $\rho^{\rm sat}$ will remain linearly stable against density fluctuations until a critical nucleus with density $\rho^{\rm in}$ and volume $V^{\rm nuc}$ is formed. We assume that the formation of the critical nucleus is a sufficiently rare event that we may consider a single nucleation event in isolation and a nucleus volume, $V^{\rm nuc}$, that is much smaller than the system volume. These assumptions provide additional limits on the range of $\rho^{\rm sat}$ in which CNT is applicable.
@@
We can make use of the condition that the droplet volume is much smaller than the system volume such that $\rho^{\rm out}\approx\rho^{\rm sat}$. In this limit, we have effectively changed our thermodynamic ensemble as the nucleus is now in contact with a particle reservoir with fixed chemical potential $\mu(\rho^{\rm sat})\equiv\mu^{\rm sat}$ and thus $\rho^{\rm in}$ can be found directly from $\mu(\rho^{\rm in})=\mu^{\rm sat}$. The critical radius is then simply ${R_{C}=2\upgamma/(\mu^{\rm sat}\Delta\rho-\Delta f})$ where the difference between an arbitrary property, $a$, in the interior and exterior of the nucleus is defined as $\Delta a\equiv a^{\rm in}-a^{\rm sat}$. The free energy barrier associated with the critical nucleus can be expressed as the sum of the volumetric and interfacial contributions ${F^{\rm{\dagger}}=4\pi R_{C}^{2}\upgamma-\frac{4}{3}\pi R_{C}^{3}(\mu^{\rm sat}\Delta\rho-\Delta f)}$ or equivalently, ${F^{\rm{\dagger}}=\frac{4}{3}\pi\upgamma R_{C}^{2}}$. The probability of observing such a critical nucleus is thus expected to be $P(R_{C})\sim\text{exp}\left[F^{{\dagger}}/k_{B}T\right]$.
It is clear that in passive systems, thermodynamics alone provides a guide for determining the density, size, and statistics of the critical nucleus. Given a saturation density, Eqs. (1) are the two equations that allow for the determination of the two unknowns: $\rho^{\rm in}$ and $R_{C}$. In Sec. II.2, we derive analogous equations capable of determining $\rho^{\rm in}$ and $R_{C}$ for systems arbitrarily far from equilibrium by adopting a mechanical perspective.
@@ II.2 Mechanics of a Critical Nucleus
Let us consider the dynamical definition of a critical nucleus. When the critical nucleus is an unstable fixed point [33], nuclei smaller (larger) than $R_{C}$ will be associated with a particle flux normal to the interface such that the nuclei shrink (grow). $R_{C}$ thus represents a critical size at which the direction of this flux switches, so the system must have zero flux normal to the interface precisely when a nucleus has a radius of $R_{C}$. For a single realization of an active fluid, the stochastic noise plays a crucial role in the density flux, which we will discuss further in Sec. II.4. For the purposes of determining the properties of the critical nucleus, we can consider an ensemble of supersaturated states. We choose a spherical coordinate system with an origin coinciding with that of the nucleus. The ensemble-averaged nucleus shape will be spherical and the ensemble-averaged density flux will no longer be stochastic. We will determine the critical nucleus size by searching for states of zero radial flux.
@@
Two flux-free states are well-established: the macroscopically phase separated state that the system will eventually reach and the initial metastable homogeneous state. The critical nucleus will correspond to an additional state in which a finite-size minority phase with zero radial flux coexists with a parent phase. The condition for determining a stationary state of a density profile is given by the linear momentum balance (see Section 1 of the SM and Ref. [34]) which governs the dynamics of the density flux. On timescales where inertia can be neglected (or for all times for overdamped particle dynamics), the radial-component of the linear momentum balance is simply $\left(\bm{\nabla}\cdot\bm{\sigma}+\mathbf{b}\right)\cdot\mathbf{e}_{r}=0$, where $\bm{\sigma}$ is the stress tensor, $\mathbf{b}$ are the body forces, and $\mathbf{e}_{r}$ is the unit vector in the radial dimension. We note that exact microscopic expressions for both $\bm{\sigma}$ and $\mathbf{b}$ can be determined through an Irving-Kirkwood-Noll procedure [35, 36, 37], an approach that we will later adopt in applying this theory to active Brownian particles. In these formulations, terms often appear to be non-local, requiring integrals over all space to, for instance, compute the stress at a particular point in space arising from particle interactions. Nevertheless, these terms are often well-approximated (e.g., when interactions are short-ranged and partial spatial correlations decay quickly) by a series of local terms, the coefficients of which can be systematically determined [38]. We thus focus on systems in which $\bm{\sigma}$ and $\mathbf{b}$ can be fully expressed as local density gradient expansions. Such a focus may become limiting for application to systems with truly long-ranged forces that prevent a local expansion of stresses/body forces.
@@
Up to a divergence-free term, we may absorb the body force into a stress-like contribution such that ${\bm{\nabla}\cdot\bm{\sigma}+\mathbf{b}=\bm{\nabla}\cdot\left[\bm{\sigma}+\mathbf{F}\right]=\bm{\nabla}\cdot\bm{\Sigma}}$, where $\mathbf{F}$ is the Coulomb integral of $\mathbf{b}$ and $\bm{\Sigma}\equiv\bm{\sigma}+\mathbf{F}$ is the dynamic stress tensor. While $\bm{\sigma}$ and $\mathbf{b}$ are fully local, the Coulomb integral may result in non-local contributions to $\mathbf{F}$. The momentum balance (${\bm{\nabla}\cdot\bm{\Sigma}=\bm{0}}$) can be used to solve for the density profile upon identifying the precise form of the dynamic stress tensor. Under spherical symmetry, $\bm{\Sigma}$ has only two independent components and can be expressed as $\bm{\Sigma}=\Sigma_{rr}\mathbf{e}_{r}\mathbf{e}_{r}+\Sigma_{tt}\left(\mathbf{I}-\mathbf{e}_{r}\mathbf{e}_{r}\right)$ where $\mathbf{I}$ is the identity tensor. We can make analytical progress by expanding the dynamic stress tensor to second order in density gradients and assuming spatial isotropy in the expansion coefficients:
$\Sigma_{tt}=-\mathcal{P}(\rho)+\kappa_{1}(\rho)\left(\frac{\partial}{\partial r}+\frac{2}{r}\right)\frac{\partial\rho}{\partial r}+\kappa_{2}(\rho)\left(\frac{\partial\rho}{\partial r}\right)^{2}+\mathcal{K},$ (2b)
@@
where $-\mathcal{P}(\rho)$ is the bulk isotropic contribution to the dynamic stress tensor, $\mathcal{K}$ is the non-local term originating from the Coulomb integral of $\mathbf{b}$, and $\{\kappa_{i}\}$ are interfacial coefficients. These interfacial coefficients can be found by formally carrying out the expansion of $\bm{\sigma}$ and $\mathbf{b}$ beginning from the microscopic definition of stress and expanding the local terms that arise from the Coulomb integral of $\mathbf{b}$. The procedure for determining the precise forms of $\{\kappa_{i}\}$ and $\mathcal{K}$ from the expansion coefficients of $\bm{\sigma}$ and $\mathbf{b}$ is outlined in Section 1.7 of the SM. In particular, a form for the nonlocal $\mathcal{K}$ term is found. We note that this form is free of any terms with explicit dependence on $1/r$. In a passive system $\mathcal{P}(\rho)$ is simply the bulk pressure and $\{\kappa_{i}\}$ are given by the Korteweg expansion [39] while non-local terms can be well-approximated by local terms for single-component systems with short-ranged interactions. More generally, the bulk pressure and interfacial coefficients can contain dynamical contributions otherwise absent in passive systems [26]. We also note that under the spherically symmetric configuration and stresses assumed here, setting the radial flux to zero also implies zero flux in the angular directions. In two dimensions, it is possible for a critical nucleus with radial symmetry to develop a finite angular flux if the underlying microscopic dynamics violate spatial parity. Describing the critical nucleus in chiral active matter [40, 41], would require a different expansion of the stress tensor such as those presented in Refs. [42, 43].
@@
Given the above physical picture, we now outline our process to find the zero radial flux states associated with a finite-sized minority phase. Our spherically symmetric geometry results in the linear momentum balance simplifying to:
@@
$\Delta\mathcal{P}=\mathcal{K}\bigr{|}_{r=0}+\frac{2\upgamma_{\rm mech}}{R_{C}}+\mathcal{O}\left(\frac{1}{R_{C}^{2}}\right).$ (4a)
@@
$=-\int_{0}^{\infty}\left[\kappa_{3}(\rho)\left(\frac{\partial\rho}{\partial r}\right)^{2}+\kappa_{4}(\rho)\frac{\partial^{2}\rho}{\partial r^{2}}\right]\text{d}r,$ (4b)
@@
where the second equality utilizes our specific form of the dynamic stress [see Eqs. (2)]. Equation (4a) is a generalized Young-Laplace equation – with the equilibrium surface energy and pressure more generally replaced by the mechanical tension and dynamic pressure, respectively, and an additional contribution arising from the non-local terms. Equation (4) carries the familiar mechanical intuition of surface tension: it measures the excess stress parallel to the interface. This mechanical tension was determined to be strongly negative for phase-separated ABPs [21] and has been the subject of some controversy regarding its physical relevance. Solon et al. [30] correctly identified that the mechanical surface tension of Kirkwood and Buff controls the Laplace pressure difference of droplets consistent with a fully local (dynamic) stress. More generally, the nonlocal term $\mathcal{K}$ may also contribute to the Laplace pressure difference. Note that Eq. (4a) ignores terms higher order in $1/R_{C}$. This approximation becomes increasingly accurate as the saturation is reduced [see Fig. 2].
@@
While the generalized Young-Laplace equation provides a mechanical condition remarkably similar to its equilibrium counterpart, we still require a second condition to allow us to determine our two unknowns, $\rho^{\rm in}$ and $R_{C}$. For passive fluids, this condition is equality of chemical potential between the nucleus and the surrounding phase. The search for chemical potential-like quantities for nonequilibrium systems has been the subject of much interest in the last decade in the context of determining macroscopic binodals [19, 45, 20, 34]. These works have resulted in generalized Maxwell constructions [20, 34] that are derived from performing the appropriate integration of the mechanical criteria of coexistence [46, 34].
@@
We follow similar arguments used in constructing the nonequilibrium coexistence criteria of macroscopic binodals [46, 34]. Beginning from the mechanical balance [Eq. (3)] we perform an indefinite integral to find an integro-differential equation for $\Sigma_{rr}$ with the Laplace pressure used as a boundary condition. We then introduce the generalized Maxwell construction pseudovariable [46, 20, 34] $\mathcal{E}$ defined by the differential equation ${(\kappa_{1}+\kappa_{4})\partial^{2}\mathcal{E}/\partial\rho^{2}=(2(\kappa_{2}+\kappa_{3})-\partial(\kappa_{1}+\kappa_{4})/\partial\rho)\partial\mathcal{E}/\partial\rho}$, which is found to be $\mathcal{E}\sim 1/\rho$ for a passive system. This pseudovariable allows us to eliminate nonlinear terms (just as it does in the case of macroscopic coexistence [46, 20, 34]) by multiplying the antiderivative of Eq. (3) by $d\mathcal{E}/dr$ and integrating from $r=0$ to $r=\infty$ to arrive at:
$\int_{\mathcal{E}^{\rm in}}^{\mathcal{E}^{\rm sat}}\text{d}\mathcal{E}\left[\mathcal{P}(\rho)-\mathcal{P}(\rho^{\rm sat})\right]+$ $\int_{0}^{\infty}\left(\mathcal{E}^{\rm in}-\mathcal{E}\right)\frac{\partial\mathcal{K}}{\partial r}\text{d}r$
$=$ $\frac{2\upgamma_{\rm ost}}{R_{C}}\frac{\rho^{\rm sat}\Delta\mathcal{E}}{\Delta\rho}+\mathcal{O}\left(\frac{1}{R_{C}^{2}}\right),$ (5a)
@@
$\upgamma_{\rm ost}=\frac{\Delta\rho}{\rho^{\rm sat}\Delta\mathcal{E}}\Biggl{[}\int_{0}^{\infty}\Bigl{(}\kappa_{1}(\rho)\frac{\partial\mathcal{E}}{\partial\rho}\left(\frac{\partial\rho}{\partial r}\right)^{2}+\mathcal{G}\left(\mathcal{E}-\mathcal{E}^{\rm in}\right)\Bigr{)}\text{d}r\Biggr{]}.$ (5b)
Note that $\upgamma_{\rm ost}$ has been defined such that it has the same units as mechanical surface tension and becomes identical to $\upgamma_{\rm mech}$ in a passive system. $\upgamma_{\rm ost}$ is named for its importance in understanding the evaporation-condensation mechanism of active coarsening, which we will discuss in Sec. II.5. As $\rho^{\rm sat}$ approaches a binodal of the phase diagram, the right hand side of Eq. (5a) vanishes and we are left with one of the macroscopic coexistence criteria. Interestingly, a term depending on gradients (and therefore the interface) remains. At equilibrium, the details of the interface do not modify the coexistence criteria as interfacial effects are generally sub-extensive in the free energy. However, outside of equilibrium no such guarantee exists. This interfacial term vanishes as the derivative of $\mathcal{K}$ approaches zero, and one may refer to the expansion of $\mathcal{K}$ in the SM for a precise expression of this limit.
@@
The connection between Eqs. (4a) and (5a) to the thermodynamic criteria shown in Eqs. (1) can be made explicit. As described by Solon et al., the generalized Maxwell constructions on the pressure can also be expressed as a difference in chemical potential-like quantities between coexisting phases [30]. It was subsequently shown that this potential satisfies a generalized Gibbs-Duhem relation [47], which, for the single-component system under consideration, is expressed as $\text{d}g=\mathcal{E}\text{d}\mathcal{P}$ where $g$ is the generalized chemical potential. The generalized Gibbs-Duhem relation allows Eq. (5a) to be expressed as a difference of chemical potentials with the nonequilibrium critical nucleus criteria now taking the form:
@@
$\Delta g=\mathcal{E}^{\rm in}\mathcal{K}\bigr{|}_{r=0}$ $+\int_{0}^{\infty}\left(2\mathcal{E}^{\rm in}-\mathcal{E}\right)\frac{\partial\mathcal{K}}{\partial r}\text{d}r$
$+\frac{2}{R_{C}}\left[\mathcal{E}^{\rm in}\upgamma_{\rm mech}+\frac{\rho^{\rm sat}\Delta\mathcal{E}}{\Delta\rho}\upgamma_{\rm ost}\right],$ (6a)
$\Delta\mathcal{P}=\mathcal{K}\bigr{|}_{r=0}+\frac{2\upgamma_{\rm mech}}{R_{C}}.$ (6b)
@@
Additionally, by defining a pseudopotential ${\Phi\equiv\int\mathcal{P}\text{d}\mathcal{E}}$, Eq. (5a) can be rearranged to express the critical radius as ${R_{C}=2\upgamma_{\rm ost}\rho^{\rm sat}\Delta\mathcal{E}/\Delta\rho\left(\Delta\mathcal{E}\mathcal{P}^{\rm sat}-\Delta\Phi\right)}$. For passive systems without non-local terms, $\upgamma_{\rm mech}=\upgamma_{\rm ost}$ and $\mathcal{E}\sim 1/\rho$. One can use these relations to readily verify that $g^{\rm in}=g^{\rm sat}$ and that Eq. (5a) recovers the thermodynamically derived expression for $R_{C}$. Our mechanical criteria for the critical nucleus [Eqs. (6)] thus recover and generalize the thermodynamic criteria [Eq. (1)].
Equations (6) [or, equivalently, Eqs. (4) and (5)] are the finite-size coexistence criteria for the critical nucleus. In Ref. [23], finite-sized coexistence criteria were derived which appear to be distinct from those of the present work. Fundamentally, this is because pressure (referred to as “pseudopressure”), surface tension (referred to as “pseudotension”), and chemical potential have different physical interpretations in Ref. [23] than in this work. Indeed, Ref. [23] explicitly cautioned against interpreting pseudopressure and pseudotension in the same manner as mechanical pressure and mechanical surface tension. In Section 1.6 of the SM we formally map AMB+ to our mechanical treatment of finite-size coexistence. Despite their ambiguous physical interpretation, we find that the finite-sized coexistence criteria presented in Ref. [23] are fully consistent with those found here. The unambiguous physical interpretation of the mechanical quantities used in our work represents an advantage of a mechanical perspective which also has a clear connection to microscopics through the Irving-Kirkwood-Noll procedure [35, 36, 37].